# Spurious Privacy Leakage in Neural Networks

### v1

@@ 4 Privacy risks of robust methods
Several methods have been proposed to mitigate spurious correlations (see Section 2). Counterintuitively, we find that mitigating the impact of spurious features does not reduce spurious privacy leakage. Although these methods improve utility fairness, group privacy disparity persists due to memorization (Zhang et al., 2021). Moreover, we demonstrate the practical limitations of differential privacy with spurious correlated data.
@@ 4.1 Privacy risks of spurious robust methods
Table 1: Comparing the privacy leakage of spurious robust methods per group. Although these methods improve the worst-group accuracy, robust training DRO and DFR do not consistently mitigate the attack across datasets. *Waterbirds is evaluated at ${\approx}3\%$ FPR due to the limited samples in the spurious groups (see Table 6 for the complete results). The spurious groups are highlighted.
@@
Experiment setup. For each dataset and training method, we train the shadow models by following the same LiRA setup as in Section 3. We ensure that models across different training methods use the same subset of data by fixing the random seeds. For privacy evaluation, we train non-overfit models by monitoring the difference between the train-val losses (Yeom et al., 2018).
@@
First, we apply spurious robust training methods to reduce overfitting of the spurious groups (see Appendix Table 5). As expected, DRO and DFR significantly reduce the difference between train-test WGA by mitigating the influence of spurious features on all five datasets (except for DRO on FMoW). We then run LiRA using ERM trained shadow models and use ERM, DRO, and DFR models as targets. Table 1 reports the privacy attack success rate for each dataset, group, and robust training method.
Figure 4: Feature similarity between ERM and DRO trained models. As a baseline, we measure feature similarity between models trained on different datasets (Waterbirds and CelebA), which results near zero similarity. Models trained with the same method (ERM or DRO) exhibit the highest alignment in feature representations. Interestingly, the similarity between ERM and DRO trained models is nearly as high, confirming that they learn similar representations despite different training objectives. Results are averaged over 5 random seeds.
@@
We observe that, across all datasets and for both spurious and non-spurious groups, the privacy vulnerability remain almost unchanged for ERM, DRO, and DFR. This indicates that spurious privacy leakage persists despite the mitigation of spurious correlations. While this result may seem surprising, overfitting is not a necessary condition for MIA to succeed (Yeom et al., 2018), which explains the ineffectiveness of spurious robust training in reducing privacy disparity. We next present an alternative perspective on privacy disparity through the lens of memorization.
@@
Memorization of spurious robust methods. We analyze spurious privacy leakage through the lens of data memorization. Memorization occurs when a model captures instance-specific features rather than learning generalizable patterns (Feldman, 2020; Feldman & Zhang, 2020). This phenomenon is tied to the success of MIA, especially for LiRA, which estimates membership by approximating the degree of memorization (Carlini et al., 2022) (see Section B.3). As shown in Figure 3, spurious groups exhibit higher memorization scores and are thus more vulnerable to LiRA than non-spurious groups. Importantly, spurious robust methods fail to reduce this memorization. ERM and DFR have nearly identical memorization score distributions across both spurious and non-spurious groups. Since DFR retrains only the final layer, it cannot address memorization which is a phenomenon distributed across multiple layers of the network (Maini et al., 2023). Therefore, DFR does not address spurious privacy leakage either. DRO shows a similar memorization level as ERM for non-spurious groups, but exhibits even higher memorization for spurious groups. Using the CKA similarity metric, we find that DRO and ERM learn highly similar feature representations (Section 3.2). Figure 4 shows that the CKA scores between embeddings from ERM and DRO are comparable to those between models trained with the same method (ERM–ERM, DRO–DRO). This aligns with findings from Izmailov et al. (2022), who showed that applying DFR on top of a DRO trained model offers no additional utility gains.
@@ 4.2 Differential privacy for spurious correlations
Differential Privacy (DP) provides formal guarantees against membership inference attacks (Dwork, 2006) (Definition D.1). For neural networks, DP-SGD (Abadi et al., 2016) enforces these guarantees by modifying the standard SGD optimizer. We audit models trained with DP using a practical evaluation, LiRA.
Experiment setup. We train target models using opacus (Yousefpour et al., 2021) following the guidelines reporte in (De et al., 2022). We use a large batch size of 1024 for Waterbirds and CelebA. The model selection uses a grid search with lr in [1, 1e-1, 1e-2, 1e-3], privacy guarantee $\epsilon$ in [1, 2, 8, 32, 128], and failure probability $\delta$ = 1e-4 for Waterbirds and $\delta$ = 1e-5 for CelebA. As in De et al. (2022); Panda et al. (2024), we do not use weight decay and the cosine scheduler due to optimization instability. Each model is trained up to 100 epochs for both Waterbirds and CelebA by selecting the best validation WGA checkpoint. For CelebA, we use a ConvNeXt-T (Liu et al., 2022b) pretrained on ImageNet-1k. For Waterbirds, we train from scratch due to dataset overlap with ImageNet-1k, which would invalidate privacy guarantees. [^1] The Waterbirds dataset combines Caltech-UCSD Birds and Places datasets; the former overlaps with ImageNet-1k (Tramèr et al., 2024) (see https://www.vision.caltech.edu/datasets/cub_200_2011/). To the best of our knowledge, CelebA has no such overlap. The privacy attack uses LiRA with 32 ERM shadow models and we report the privacy budget as $(\epsilon,\delta)$.
Figure 5 (left) presents the results of large batch trained target models across four metrics, the average and worst-group utility and privacy metrics as a function of the privacy budget $\epsilon$. Prior works by Bagdasaryan et al. (2019) shows that DP training with small batch size increases the group utility disparity, harming the worst-group utility. However, consistent with Panda et al. (2024), we observe that DP training with large batch size can help to mitigate this disparity. For example, on CelebA, a tight budget of $\epsilon=1$ has the lowest accuracy, but a relaxed $\epsilon$ improves the worst-group accuracy. Note that this trend does not hold for models trained from scratch on Waterbirds, which lacks a “good” weight initialization.
@@
For privacy mitigation, we observe a utility–privacy trade-off in Figure 5 (right). On CelebA, while average privacy remains relatively stable, stricter $\epsilon$ values can improve privacy protection for the worst group. In contrast, Waterbirds shows no significant change, with both the average and worst-class TPR remaining consistent across $\epsilon$. In summary, while differential privacy can improve utility fairness and privacy protection, its effectiveness depends on the presence of a DP suitable pretrained model (Tramèr et al., 2024) to also mitigate spurious correlation.
@@ 5 Architecture influence on spurious correlations and privacy
Prior privacy works mostly focused on settings with ResNet-like architecture. However, modern architectures have different components that impact feature learning (e.g. masking from He et al. (2022) or attention from Dosovitskiy et al. (2021)). We evaluate models that are sufficiently “diverse”, comparing models from different families (convolution and transformers), with different pretraining strategies (supervised and self-supervised), and released at different times (e.g. ResNet and its successor ConvNext). We provide counterexamples to revisit prior research findings. First, vision transformers are not more spurious robust than convolution models under a fair evaluation setup. Second, the best shadow architecture does not always match with the target architecture.
Experiment setup. All the models used are pretrained using the state-of-the-art recipe on the ImageNet1K dataset from the timm library. We compare ResNet50 (He et al., 2016), BiT-S (Kolesnikov et al., 2020), CNext-T (Liu et al., 2022b), CNextV2-T (Woo et al., 2023), ViT (Dosovitskiy et al., 2021), Swin-T (Liu et al., 2021b), DeiT-S (Touvron et al., 2022), and Hiera (Ryali et al., 2023) on Waterbirds. To ensure a fair comparison, all the models have a similar number of parameters, ranging from 20M to 30M. We train 16 shadow models for each architecture as in Section 3 while monitoring the train-validation loss to prevent overfitting. Appendix Table 8 reports the details about the grid search and summarizes the best hyperparameters. The results are averaged over 16 seeds.
Are ViTs more spurious robust than CNNs? Ghosal & Li (2024) claim that vision transformers are more robust than convolutional models. We revisit this statement, showing that under the same grid search, both architectures perform comparably. Table 2 shows that the best convolutional and transformer based models (Swin and CNextV2) have similar WGA of ${\approx 52\%}$. Moreover, our grid search on the architecture DeiT-S/16 achieves +4.7% absolute improvement in WGA over the value reported by Ghosal & Li (2024) (Table 4), raising concerns about their evaluation protocol. We also note that they unfairly compare the pretrained convolution model BiT with ViT/DeiT, since the latter is pretrained on ImageNet1K with a complex optimization recipe (RandAug, Mixup, CutMix, LabelSmoothing, StochasticDepth, LayerScale) while BiT is not. Our benchmark addresses these issues by comparing a broader set of architectures under a fixed compute budget, ensuring fair evaluation. Additional consistent results are reported for the CelebA dataset in the Appendix Table 9.
@@
Pretraining recipe and architecture matter for privacy auditing. Prior works audit the privacy of different architectures on well-balanced datasets such as CIFAR or ImageNet (Shokri et al., 2017; Carlini et al., 2022; Liu et al., 2022a). Our setup compares state-of-the-art architectures on non-balanced datasets Waterbirds, showing that architecture choice can matter for privacy auditing. Using LiRA, we run the attack through all the permutations of shadow/target architecture pairs, resulting in 64 different configurations. Figure 6 shows that when we fix the shadow architecture and vary the target, there is no particular architecture that is more resistant to others. On contrary, when fixing the target architecture while varying the shadow, we observe that some architectures are consistently achieve the strongest attack across all the targets. Moreover, despite sharing the exact same architecture with ViT but differing in optimization recipe, DeiT is consistently ranked as the lowest, suggesting that the pretraining optimization can greatly influence the attack success of shadow models. Lastly, prior works reported that the most successful attack is when the shadow matches the target architecture (Carlini et al., 2022; Liu et al., 2022a). In our setup with spurious data, we do not observe the same trend. In Figure 6, the black-outlined markers represent the match between shadow and target architectures, which corresponds to a suboptimal attack. Additional consistent results are reported for the CelebA dataset in the Appendix Figure 9.
@@ 6 Conclusion
Our findings highlight critical privacy concerns when training neural networks on real-world datasets with spurious correlations. The existence of spurious privacy leakage can make spurious data significantly more vulnerable to privacy attacks than non-spurious data. Leveraging this information, an adversary can craft better tailored attacks against specific demographic groups. To detect and prevent this, the auditor can go beyond aggregate metrics and include a fine-grained, group-level analysis of both performance and privacy fairness. This is particularly important for datasets with simpler task objective (Section 3.2).
We further show that existing strategies are not sufficient. For example, the most promising approaches of spurious robust methods fail to address spurious privacy leakage. Our analysis attributes the failure to high feature similarity between robust and standard training models, demonstrating that the underlying memorization behavior remains unchanged. Moreover, alternative spurious methods (Section 2) effectively reduce to variants of either DRO or DFR relying on group resampling or sample reweighting, and thus, they are likely to show similar limitations. Although differential privacy offers formal guarantee, a tight privacy budget comes at the expense of the worst-group utility, which is not practical. These findings open several promising directions for future research such as developing private spurious robust methods or better auditing tools to detect group disparicy without accessing group labels. Lastly, we revisit prior work on the relationship between architecture, spurious correlations, and privacy, providing insights that revisit and complement existing research.
Impact. Our findings impacts the machine learning communities concerned with bias, fairness, and security. Understanding the connection between spurious correlations and privacy helps to identify the risks for data-sensitive applications. In particular, we suggest practitioners working in auditing to carefully assess privacy disparities and avoid the use of average metrics for spurious correlated data.
Limitations. Our results are based on a state-of-the-art attack-based evaluation rather than analytical guarantees. While our approach is more practical and provides empirical evidence, it is limited to the choice of our experiment settings, which include five real-world datasets, eight model architectures, and two robust training methods.
Reproducibility. We have made our code publicly available through an anonymized repository (Section 1). Details for each experiment setup are presented in the respective sections, along with the corresponding grid search and the best hyperparameters in the appendix (see Tables 3 and 8). The spurious datasets are presented in the Appendix A.
@@ Acknowledgments
CZ is funded by the research project R-STR-8019-00-B from the University of Luxembourg. The experiments are supported by the ULHPC facilities from the University of Luxembourg.  [^2] https://hpc-docs.uni.lu/

### v2

@@ 4 Privacy risks of robust methods
Several methods have been proposed to mitigate spurious correlations (see Section˜2). Counterintuitively, we find that mitigating the impact of spurious features does not reduce spurious privacy leakage. Although these methods improve utility fairness, group privacy disparity persists due to memorization (Zhang et al., 2021). Moreover, we demonstrate the practical limitations of differential privacy with spurious correlated data.
@@ 4.1 Privacy risks of spurious robust methods
Table 1: Comparing the privacy leakage of spurious robust methods per group. Although these methods improve the worst-group accuracy, DRO and DFR do not consistently mitigate the attack across datasets. *Waterbirds is evaluated at ${\approx}3\%$ FPR due to the limited samples in the spurious groups (see Table˜6 for the complete results). Bolded values represent the best training method for privacy mitigation. The spurious groups are highlighted.
@@
Experiment setup. For each dataset and training method, we train the shadow models by following the same LiRA setup as in Section˜3. We ensure that models across different training methods use the same subset of data by fixing the random seeds. For privacy evaluation, we train non-overfit models by monitoring the difference between the train-val losses (Yeom et al., 2018).
@@
First, we apply spurious robust training methods to reduce overfitting of the spurious groups (see Appendix Table˜5). As expected, DRO and DFR significantly reduce the difference between train-test WGA by mitigating the influence of spurious features on all five datasets (except for DRO on FMoW). We then run LiRA using ERM trained shadow models and use ERM, DRO, and DFR models as targets. Table˜1 reports the privacy attack success rate for each dataset, group, and robust training method. We observe that, across all datasets and for both spurious and non-spurious groups, the privacy vulnerability remain almost unchanged for ERM, DRO, and DFR. This indicates that spurious privacy leakage persists despite the mitigation of spurious correlations. While this result may seem surprising, overfitting is not a necessary condition for MIA to succeed (Yeom et al., 2018), which explains the ineffectiveness of spurious robust training in reducing privacy disparity. We next present an alternative perspective on privacy disparity through the lens of memorization.
Memorization of spurious robust methods. We analyze spurious privacy leakage through the lens of data memorization. Memorization occurs when a model captures instance-specific features rather than learning generalizable patterns (Feldman, 2020; Feldman & Zhang, 2020). This phenomenon is tied to the success of MIA, especially for LiRA, which estimates membership by approximating the degree of memorization (Carlini et al., 2022) (see Section˜B.3).
Figure 4: Feature similarity of different training objectives using linear-CKA. For the baseline, we have near-zero similarity between models trained on different datasets (Waterbirds and CelebA). A pair of models trained with different training methods (ERM, DRO) on the same dataset exhibits similar similarity as those trained using the same method, (ERM, ERM) or (DRO, DRO). Results are averaged over 5 random seeds.
@@
As shown in Figure˜3, spurious groups exhibit higher memorization scores and therefore are also more vulnerable to LiRA than non-spurious groups. Importantly, spurious robust training fails to reduce this memorization. The methods ERM and DFR have nearly identical memorization score distributions across both spurious and non-spurious groups. Since DFR re-trains only the final layer, it cannot address memorization which is a phenomenon distributed across multiple layers of the network, as demonstrated by Maini et al. (2023). Therefore, DFR does not address spurious privacy leakage either. DRO shows a similar memorization level as ERM for non-spurious groups, but exhibits even higher memorization for spurious groups. Using the CKA similarity metric, we investigate the embedding similarity of models trained with ERM and DRO. We find that the learned embeddings have highly similar feature representations (Section˜3.2). In particular, Figure˜4 shows that the CKA scores between embeddings from ERM and DRO are comparable to those between models trained with the same spurious method (ERM–ERM or DRO–DRO). This aligns with findings from Izmailov et al. (2022), who showed that applying DFR on top of a DRO trained model offers no additional utility gains.
@@ 4.2 Differential privacy for spurious correlations
Differential Privacy (DP) provides formal guarantees against membership inference attacks (Dwork, 2006) (Definition˜D.1). For neural networks, DP-SGD (Abadi et al., 2016) enforces these guarantees by modifying the standard SGD optimizer. We audit models trained with DP using a practical evaluation, LiRA.
Experiment setup. We train target models using opacus (Yousefpour et al., 2021) following the guidelines from De et al. (2022). We use a large batch size of 1024 for Waterbirds and CelebA. The model selection uses a grid search with lr in [1, 1e-1, 1e-2, 1e-3], privacy guarantee $\epsilon$ in [1, 2, 8, 32, 128], and failure probability $\delta$ = 1e-4 for Waterbirds and $\delta$ = 1e-5 for CelebA. As in De et al. (2022); Panda et al. (2024), we do not use weight decay and the cosine scheduler due to optimization instability. Each model is trained up to 100 epochs for both Waterbirds and CelebA by selecting the best validation WGA checkpoint. For CelebA, we use a ConvNeXt-T (Liu et al., 2022b) pretrained on ImageNet-1k. For Waterbirds, we train from scratch due to dataset overlap with ImageNet-1k, which would invalidate privacy guarantees. [^1] The Waterbirds dataset combines Caltech-UCSD Birds and Places datasets; the former overlaps with ImageNet-1k (Tramèr et al., 2024) (see https://www.vision.caltech.edu/datasets/cub_200_2011/). To the best of our knowledge, CelebA has no such overlap. The privacy attack uses LiRA with 32 ERM shadow models and we report the privacy budget as $(\epsilon,\delta)$.
Figure˜5 (left) presents the results of large batch trained target models across four metrics, the average and worst-group utility and privacy metrics as a function of the privacy budget $\epsilon$. Prior works by Bagdasaryan et al. (2019) shows that DP training with small batch size increases the group utility disparity, harming the worst-group utility. However, consistent with Panda et al. (2024), we observe that DP training with large batch size can help to mitigate this disparity. For example, on CelebA, a tight budget of $\epsilon=1$ has the lowest accuracy, but a relaxed $\epsilon$ improves the worst-group accuracy. Note that this trend does not hold for models trained from scratch on Waterbirds, which lacks a “good” weight initialization.
@@
For privacy mitigation, we observe a utility–privacy trade-off in Figure˜5 (right). On CelebA, while average privacy remains relatively stable, stricter $\epsilon$ values can improve privacy protection for the worst group. In contrast, Waterbirds shows no significant change, with both the average and worst-class TPR remaining consistent across $\epsilon$. In summary, while differential privacy can improve utility fairness and privacy protection, its effectiveness depends on the presence of a DP suitable pretrained model (Tramèr et al., 2024) to also mitigate spurious correlation.
@@ 5 Architecture influence on spurious correlations and privacy
Prior privacy works mostly focused on settings with ResNet-like architecture. However, modern architectures have different components that impact feature learning (e.g. masking from He et al. (2022) or attention from Dosovitskiy et al. (2021)). We evaluate models that are sufficiently “diverse”, comparing models from different families (convolution and transformers), with different pretraining strategies (supervised and self-supervised), and released at different times (e.g. ResNet and its successor ConvNext). We provide counterexamples that revisit two prior research findings. First, vision transformers are not more robust than convolution models under a fair spurious evaluation setup. Second, the best shadow architecture does not always match the target architecture.
Experiment setup. All the models used are pretrained using the state-of-the-art recipe on the ImageNet1K dataset from the timm library. We compare ResNet50 (He et al., 2016), BiT-S (Kolesnikov et al., 2020), CNext-T (Liu et al., 2022b), CNextV2-T (Woo et al., 2023), ViT (Dosovitskiy et al., 2021), Swin-T (Liu et al., 2021b), DeiT-S (Touvron et al., 2022), and Hiera (Ryali et al., 2023) on Waterbirds. To ensure a fair comparison, all the models have a similar number of parameters, ranging from 20M to 30M. We train 16 shadow models for each architecture as in Section˜3 while monitoring the train-validation loss to prevent overfitting. Appendix Table˜8 reports the details about the grid search and summarizes the best hyperparameters. The results are averaged over 16 seeds.
Are ViTs more spurious robust than CNNs? Ghosal & Li (2024) claim that vision transformers are more robust than convolutional models. We revisit this statement, showing that under the same grid search, both architectures perform comparably. Table˜2 shows that the best convolutional and transformer based models (Swin and CNextV2) have similar WGA of $\approx 52\%$}. Moreover, our grid search on the architecture DeiT-S/16 achieves +4.7% absolute improvement in WGA over the value reported by Ghosal & Li (2024) (Table 4), raising concerns about their evaluation protocol. We also note that they unfairly compare the pretrained convolution model BiT with ViT/DeiT, since the latter is pretrained on ImageNet1K with a complex optimization recipe (RandAug, Mixup, CutMix, LabelSmoothing, StochasticDepth, LayerScale) while BiT is not. Our benchmark addresses these issues by comparing a broader set of architectures under a fixed compute budget, ensuring fair evaluation. Additional consistent results are reported for the CelebA dataset in the Appendix Table˜9.
@@
Pretraining recipe and architecture matter for privacy auditing. Prior works audit the privacy of different architectures on well-balanced datasets such as CIFAR or ImageNet (Shokri et al., 2017; Carlini et al., 2022; Liu et al., 2022a). Our setup compares state-of-the-art architectures on non-balanced datasets Waterbirds, showing that architecture choice can matter for privacy auditing. Using LiRA, we run the attack through all the permutations of shadow/target architecture pairs, resulting in 64 different configurations. Figure˜6 shows that when we fix the shadow architecture and vary the target, there is no particular architecture that is more resistant than others. On the contrary, when fixing the target architecture while varying the shadow, we observe that some architectures consistently achieve the strongest attack across all the targets. Moreover, despite sharing the exact same architecture with ViT but differing in optimization recipe, DeiT is consistently ranked as the lowest, suggesting that the pretraining optimization can greatly influence the attack success of shadow models. Lastly, prior works reported that the most successful attack is when the shadow matches the target architecture (Carlini et al., 2022; Liu et al., 2022a). In our setup with spurious data, we do not observe the same trend. In Figure˜6, the black-outlined markers represent the match between shadow and target architectures, which corresponds to a suboptimal attack. Additional consistent results are reported for the CelebA dataset in the Appendix Figure˜10.
@@
Lastly, further results in the Appendix (Figures˜9 and 11) show that all the evaluated architectures exhibit spurious privacy leakage. Across architectures (e.g. Resnet, Convnext, ViT, and Swin), the difference between the most and the least spurious group is approximately one order of magnitude under TPR at 0.1% FPR (see Figure˜9 groups 2/3 for Waterbirds and Figure˜11 groups 3/1 for CelebA). However, note that the privacy leakage across all the different target architectures is at a similar level, suggesting that architecture choice alone plays a small role in mitigating spurious privacy leakage.
@@ 6 Conclusion
Our findings highlight critical privacy concerns when training neural networks on real-world datasets with spurious correlations. The existence of spurious privacy leakage makes spurious data significantly more vulnerable to privacy attacks than non-spurious data. Using this information, an adversary can craft better-tailored attacks against specific demographic groups. To detect and prevent this, the auditor can go beyond aggregate metrics and include a fine-grained group-level analysis of both performance and privacy fairness. This is particularly important for datasets with simpler task objectives (Section˜3.2).
We further show that existing strategies are not sufficient. For example, the most promising approaches of spurious robust methods fail to address spurious privacy leakage. Our analysis attributes the failure to high feature similarity between robust and standard training models, demonstrating that the underlying memorization behavior remains unchanged. Moreover, alternative spurious methods (Section˜2) effectively reduce to variants of DRO or DFR relying on group resampling or sample reweighting and therefore are likely to show similar limitations. Although differential privacy offers a formal guarantee, a tight privacy budget comes at the expense of the worst-group utility, which is not practical.
Our results open promising directions for future research. On the attack side, privacy disparity enables adversaries to achieve higher MIA success by targeting only subgroups of data. For example, an adversary can inject poisoned samples (tramer2022truth) to strengthen a spurious correlation, allowing the attacker to target a subgroup with a higher success rate. Another possible direction is to improve the auditing of spurious data in real world without explicit spurious attribute labels. For example, label-free spurious robust methods (Yang et al., 2024) can be used to detect spurious correlations and perform auditing of both performance and privacy fairness, mitigating undesired deployment.
Limitations. Our results are based on a state-of-the-art attack-based evaluation rather than analytical guarantees. Although our approach is more practical and provides empirical evidence, it is limited to the choice of our experiment settings, which include five real-world datasets, eight model architectures, and two robust training methods.
Reproducibility. We have made our code publicly available through an anonymized repository (Section˜1). Details for each experiment setup are presented in the respective sections, along with the corresponding grid search and the best hyperparameters in the Appendix (see Tables˜3 and 8). The spurious data sets are presented in Appendix˜A. The code is released at https://github.com/orientino/spurious-mia.
@@ Acknowledgments
CZ is funded by the research project R-STR-8019-00-B from the University of Luxembourg. The experiments were supported by the ULHPC facilities from the University of Luxembourg.  [^2] https://hpc-docs.uni.lu/